# MetaMetrics: Calibrating Metrics For Generation Tasks Using Human Preferences

### v1

@@ 1 Introduction
  • We demonstrate that MetaMetrics is adaptable to a wide range of requirements, including performance optimization and metric efficiency through a comprehensive benchmarking of diverse metrics across a range of language and vision tasks Section 4. We also show that MetaMetrics can be used effectively as a reward model. [^1] MetaMetrics outperforms the state-of-the-art reward model reported at RewardBench leaderboard on October 1st, 2024. This adaptability ensures that MetaMetrics can be calibrated to suit various human preference aspects.
@@ 5 Results
MetaMetrics-RM
GP 94.5 98.3 86.4 90.8 98.0
XGBoost 93.5 98.3 84.2 89.7 97.1
XGBoost (Iterative Best) 93.5 98.3 84.2 89.7 97.1
Table 4: Accuracy results in percentage for Reward-Model-As-A-Metric. More detailed results for all metrics and their variants can be found in Table 13 in the Appendix. Bolded and underlined values indicate the best and second best performance, respectively. Our MetaMetrics-RM performs the best or comparable on 2 of the 4 categories in RewardBench.
@@
Table 4 highlights the performance of MetaMetrics-RM, which achieves state-of-the-art results [^5] We use the RewardBench leaderboard from October 1, 2024, as the basis for comparison. for GP and competes effectively with the top seven metrics, including the advanced 70B Nemotron Reward model. These results demonstrate that our proposed metric is both efficient and effective, showcasing its robustness across diverse categories of tasks, datasets, and evaluation settings. Overall, MetaMetrics-RM represents a significant advancement in reward metric evaluation, offering high performance and broad applicability for researchers and practitioners alike.
@@ 6.2 Efficiency
Table 4 highlights the compute efficiency and superior performance of MetaMetrics-RM. MetaMetrics-RM, utilizing GP, combines GRM-Gemma-2B, Internlm2-7B-Reward, Skywork-Reward-Llama-3.1-8B, Internlm2-1.8B-Reward, and GRM-Llama3-8B totaling approximately 27B parameters. Despite this relatively small compute memory footprint, it outperforms models more than 2$\times$ its size, achieving 0.4% higher overall accuracy than the current state-of-the-art. This demonstrates that MetaMetrics-RM offers state-of-the-art accuracy using significantly more cost-effective models, making it an ideal choice for resource-constrained applications while delivering even better performance.
MetaMetrics also excels at selecting a sparse yet highly effective set of metrics in scenarios with a vast number of potential evaluation metrics. As depicted in Figure 4, both GP and XGBoost are capable of identifying small subset of key metrics that correlate strongly with human preferences, reducing the need to evaluate a wide range of metrics. This selective approach identifies key metrics to a specified task while minimizing computational overhead without compromising performance. In other words, MetaMetrics is well-suited in low-resource settings while still providing better alignment with human preferences across multiple tasks.
@@ C.2.1 Bayesian Optimization
Table 7 describes the hyper-parameter settings that we use for our experiments.
@@ D.1 RewardBench Leaderboard
Figure 5: RewardBench Leaderboard Screenshot (Updated on October 1st, 2024). Our best model outperforms the state-of-the-art model.
@@ D.6 Reward Model Scoring
GP 94.5 98.3 86.4 90.8 98.0
GP (Top 3) 94.0 95.5 87.7 91.6 96.9
XGBoost 93.5 98.3 84.2 89.7 97.1
XGBoost (Iterative Best) 93.5 98.3 84.2 89.7 97.1
XGBoost (Top 3) 93.1 96.9 85.8 91.2 95.5
@@ E.2 Machine Translation
There are 2 setups for MT, reference-based and reference-free. Both setups’ resulting weights or feature importance are reported in 7

### v2

@@ 1 Introduction
  • We demonstrate that MetaMetrics is adaptable to a wide range of requirements, including performance optimization and metric efficiency through a comprehensive benchmarking of diverse metrics across a range of language and vision tasks Section 4. We also show that MetaMetrics can be used effectively as a reward model. [^1] MetaMetrics ranks 4th out of 144 on the overall leaderboard, utilizing the fewest parameters, and secures 2nd place on the detailed leaderboard at RewardBench. The leaderboard reflects data as of October 6th, 2024. This adaptability ensures that MetaMetrics can be calibrated to suit various human preference aspects.
@@ 5 Results
GRM-Llama3-8B 91.5 95.5 86.2 90.8 93.6
ArmoRM-Llama3-8B-v0.1 90.4 96.9 76.8 90.5 97.3
URM-LLaMa-3-8B 89.9 96.9 78.7 88.2 95.7
MetaMetrics-RM (GP) 93.4 98.3 86.4 90.8 98.2
MetaMetrics-RM-XGBoost
   Iterative Best 92.9 95.8 88.4 91.8 95.8
Table 4: Accuracy results in percentages for Reward-Model-As-A-Metric. More detailed results for all metrics and their variants can be found in Table 13 in the Appendix. Values that are bolded indicate the best performance, while those that are underlined represent the second-best performance. Our MetaMetrics-RM achieves the best or comparable results in 2 out of the 4 categories in RewardBench.
@@
Table 4 highlights the performance of MetaMetrics-RM, which achieves state-of-the-art results [^5] We use the RewardBench leaderboard from October 6th, 2024, as the basis for comparison. for GP and competes effectively with the top seven metrics, including the advanced 70B Nemotron Reward model. These results demonstrate that our proposed metric is both efficient and effective, showcasing its robustness across diverse categories of tasks, datasets, and evaluation settings. Overall, MetaMetrics-RM represents a significant advancement in reward metric evaluation, offering high performance and broad applicability for researchers and practitioners alike.
@@ 6.2 Efficiency
Improved Performance with Fewer Parameters.
Table 4 highlights the compute efficiency and superior performance of MetaMetrics-RM. MetaMetrics-RM, utilizing GP, combines GRM-Gemma-2B, Internlm2-7B-Reward, Skywork-Reward-Llama-3.1-8B, Internlm2-1.8B-Reward, and GRM-Llama3-8B totaling approximately 27B parameters. Despite this relatively small compute memory footprint, it outperforms models more than 2$\times$ its size. This demonstrates that MetaMetrics-RM offers state-of-the-art accuracy using significantly more cost-effective models, making it an ideal choice for resource-constrained applications while delivering even better performance.
High Parallelization.
MetaMetrics is designed to be embarrassingly parallel, enabling each metric to be executed independently without requiring inter-model communication, which results in exceptional efficiency. Unlike large models with 27 billion parameters that often face bottlenecks, our approach supports the development of highly effective models without such time constraints. Furthermore, MetaMetrics excels at selecting a sparse yet impactful set of metrics from a vast pool of potential evaluation metrics. As shown in Figure 4, both GP and XGBoost can identify a small subset of key metrics that strongly correlate with human preferences, thereby reducing the need to evaluate a wide range of metrics. This selective approach identifies essential metrics for specific tasks while minimizing computational overhead without compromising performance. In essence, MetaMetrics is ideally suited for low-resource settings, providing superior alignment with human preferences across multiple tasks.
@@ C.2.1 Bayesian Optimization
Table 7 describes the hyper-parameter settings that we use for our experiments. For the Bayesian optimization, we run GP with a Matern kernel Williams & Rasmussen (2006), a generalization of the RBF kernel, using $\nu=2.5$.
@@ D.1 RewardBench Leaderboard
Figure 5: RewardBench Leaderboard Screenshot (Updated on October 6th, 2024).
Figure 6: RewardBench Detailed Leaderboard Screenshot (Updated on October 6th, 2024).
@@ D.6 Reward Model Scoring
GP 93.4 98.3 86.4 90.8 98.2
GP (Top 3) 93.0 95.5 87.7 91.6 97.0
XGBoost 92.8 96.1 88.4 90.9 95.8
XGBoost (Iterative Best) 92.9 95.8 88.4 91.8 95.8
XGBoost (Top 3) 93.0 86.0 90.3 93.0 95.5
@@ E.2 Machine Translation
There are 2 setups for MT, reference-based and reference-free. Both setups’ resulting weights or feature importance are reported in 8